OpenStudy (xapproachesinfinity):

check out this question: find \[f'(\dfrac{\pi}{2})\] \[\large f(x)=\int_{0}^{g(x)} \frac{1}{\sqrt{1+t^3}}dt \], where \[\large g(x)=\int_{0}^{\cos x} [1+\sin (t^2)]dt\] see if you can find -1 like I found or perhaps i made a mistake there...

OpenStudy (xapproachesinfinity):

i can't believe i posted this in feedback section lol that hilarious :)

OpenStudy (freckles):

\[f'(x)=\frac{g'(x)}{\sqrt{1+(g(x))^3}} \\ g'(x)=-\sin(x) \cdot (1+\sin((\cos(x))^2) \\ g'(\frac{\pi}{2})=-1(1+\sin((0))=-1 \\ f'(\frac{\pi}{2})=\frac{-1}{\sqrt{1+(g(\frac{\pi}{2}))^3}} \\ g(\frac{\pi}{2})=\int\limits_{0}^{0}(blah) dt=0 \\ f'(\frac{\pi}{2})=-1 \]

OpenStudy (freckles):

yep i seem to get -1 too

OpenStudy (freckles):

also this in is math

OpenStudy (freckles):

according to my screen

OpenStudy (xapproachesinfinity):

oh great :)

OpenStudy (xapproachesinfinity):

i posted first in feedback hehe

OpenStudy (freckles):

oh

OpenStudy (xapproachesinfinity):

i kind of did it differently, let me analyze your way ^^

OpenStudy (xapproachesinfinity):

oh its the same actually, it thought i was something odd lol

OpenStudy (xapproachesinfinity):

thanks

OpenStudy (xapproachesinfinity):

saw *

OpenStudy (freckles):

lol the blah thing?

OpenStudy (xapproachesinfinity):

yeah :)

OpenStudy (xapproachesinfinity):

there is problem in max, min i couldnt solve yet i will give it some more time and see if i can come up with something